3. Evaluate each of the following by first expressing each in terms of factorials. a) „P; b) P(8, 3) c) 54C53 d) e) P(9,0)
Answer #1

We know that, upra ne nr = > (n-oy! al ( no)! a) is P2 181 (18-2)! 18 x 17 x 14! 306 b) P(8,3)= 8P3 = ! (8-3). = 8x7x6x8! St.54C53 - 54! 53! (54-53) 54x57. 53. ! = 54 a) sula = 54! 54! 2!. (54-2)! 27 Sax53x571 9x1x5%! 1431 es p(9,0) = q po (9-0)